# Dividing Indivisible Items for the Benefit of All: It is Hard to Be Fair Without Social Awareness

### v1

@@ Abstract
1 Introduction
@@
The scenarios above might be seen as fair division with externalities (Velez 2016; Aziz et al. 2023b; Deligkas et al. 2024), where agents care not only for the items assigned to them, but also how exactly the remaining items are allocated. Nevertheless, this model is too general, so Flammini, Greco, and Varricchio (2025) have recently introduced a model that captures exactly the setting above: there is a set of goods and every agent is associated with two additive functions that define their value and social impact for each good. The goal is to allocate the goods so that the social impact is maximized while maintaining some fairness criterion. They have studied the “price of fairness”, which essentially quantifies how much worse in terms of social welfare a fair outcome is. In addition, they have formally introduced the notion of socially aware agents and which outcomes they perceive as fair. They showed that with socially aware agents, EF1 and social impact maximizing allocations always exist and can be computed in polynomial time. However, the complexity of the underlying computational problems was not examined in detail. Our goal is to extend the fairness solution concepts for this model and settle their complexity.
Our Contribution
@@
Socially-Aware Agents All Agents Socially-Aware yes NP-complete [Thm. 9] no $\alpha$-Approx. Socially-Aware yes NP-complete [Thm. 11] $\alpha<1$ Weakly Socially-Aware $\alpha=1$ NP-complete [Thm. 12] yes poly-time [Thms. 5, 6] no Binary Valuations no Valuation = Social-Impact yes poly-time [Thm. 2] yes NP-complete [Thm. 1] no Constant # of Agents no NP-complete [Thm. 1] no Bounded Valuations yes NP-complete [Thm. 3] no poly-time [Thm. 4] yes
@@
Figure 1: A basic overview of the complexity landscape of deciding whether a SIM and $\mathcal{F}$-fair allocation exist.
@@
We provide a comprehensive study of the complexity of computing fair allocations that maximize social impact, which we term SIM allocations. We investigate seven different fairness criteria, all of which are strengthenings or weakenings of EF1. For each criterion, we examine the constraints on the valuation and social impact functions that allow for tractability (see Figure 1 for an overview).
@@
Our initial set of technical results considers socially unaware agents, where we study two different dimensions of the problem and for each of them we provide complementary results between tractability and NP-hardness. Interestingly, we show that the problem behaves in the same way for every fairness criterion, albeit we need slight modifications in our proofs to formally establish this. First, we restrict the domain of the valuation and social impact functions. There, we show that the problem is NP-hard even when both functions are binary (Thm. 1), but it becomes tractable if for every agent both functions are binary and equal to each other (Thm. 2). Then, we restrict the number of agents. Unfortunately, the problem is NP-hard even for two agents with binary social impact functions (Thm. 3). On the other hand, we derive a pseudo-polynomial time algorithm that solves the problem for any constant number of agents (Thm. 4). Hence, the problem is tractable when we have a constant number of agents, polynomially bounded valuation functions, and unrestricted social impact functions.
@@
Then we consider the case where the agents are socially aware. Our first set of results for this setting adopts the definition of social awareness from (Flammini, Greco, and Varricchio 2025). Intuitively, we accept allocations that do not satisfy the “standard” fairness constraint, however, any bundle an agent envies (according to the chosen fairness criterion) cannot produce strictly more social impact if this agent gets it. We prove that social awareness allows for SIM and fair allocations that can be computed in polynomial time, for all fairness criteria we study (Thms. 5 and 6). This leads us to the question of whether the definition of social awareness allows for an arbitrary fairness criterion. We prove that this is not the case by showing NP-hardness if we have “fully envious” agents (Thm. 7).
@@
Finally, we ask whether we can relax social awareness and still get a fair and SIM allocation in polynomial time. Our first relaxation considers the case where not all agents are socially aware. Unfortunately, even if only one agent is socially unaware existence of SIM and fair allocations is not guaranteed and, in fact, the problem becomes NP-hard even when there are two agents (Thm. 9). On the positive side, we complement this negative result with a pseudo-polynomial time algorithm that works for any constant number of agents where some of them are socially unaware (Thm. 10). Then, we limit social awareness for each agent in two different ways. In the first case, we focus on $\alpha$-approximately socially aware agents, where $\alpha\in[0,1]$, and an agent accepts an unfair outcome only if every envied bundle produces at least $\alpha$ times more social impact under the current allocation. Strikingly, we get the same behavior as before: there is not always a solution and the problem is NP-hard for every possible $\alpha$ (Thm. 11). In the second case, we define weakly socially aware agents. Intuitively, an agent accepts an unfair allocation if the proportional gain of their valuation from an envied bundle is less than the proportional gain in social impact (Def. 11). Once again, there is no guaranteed solution and the problem is NP-hard even for two agents (Thm. 12).
Related Work
@@
Our paper lies in the area of fair division with indivisible items, where there is a flourish of results in the last 15 years (Lipton et al. 2004; Bouveret and Lang 2008; Budish 2011; Caragiannis et al. 2019); for excellent recent surveys, see (Amanatidis et al. 2023) and (Nguyen and Rothe 2023).
@@
As we have mentioned before, (Flammini, Greco, and Varricchio 2025) is the most relevant paper to ours, since we extend their model and their computational results. Having said this, both papers fall into the larger, more general model of fair division with externalities (Velez 2016; Aziz et al. 2023b; Deligkas et al. 2024). In this model, each agent gets value from the whole allocation, i.e., they value each bundle depending on the agent who got it. Hence, we can see any instance in our setting as an instance of agents with externalities. On a similar note, Bu et al. (2023) studied the model in which there is an allocator with its own view of how the items should be allocated between agents, and the goal is to be fair to both the agents and the allocator. It should be highlighted that users deriving their utility not only from their own benefit but also from the benefit of (or to) others is a well-described phenomenon in social sciences and behavioral economics; see, e.g., (Overvold 1980; Andreoni 1990; Huang et al. 2022; Thomas and Thomas 2025).
@@
A different point of view is to consider our problem as constrained fair division (Suksompong 2021). We identify two settings that are closest to ours. The first regards welfare-maximizing fair allocations. Observe that when social impact equals the valuation of each agent, fair SIM allocations coincide with welfare-maximizing allocations that are fair, which was proven to be NP-hard (Aziz et al. 2023a) in general. The second is related to orientations (Christodoulou et al. 2023; Zhou et al. 2024; Deligkas et al. 2025; Afshinmehr et al. 2025; Blažej et al. 2025). Here, every agent is associated with a subset of items they are interested in. The goal is to find an allocation where every agent gets items only from their subset. In our case, agents are only allowed to get items for which they are social impact maximizers. The crucial difference though is that in our case, agents value the remaining items as well.
2 Preliminaries
@@
We assume a set M{M} of m{m} items which need to be partitioned between a set N{{N}} of n{{n}} agents. Each agent $i\in{{N}}$ is associated with a valuation function ${v}_{i}\colon 2^{M}\to\mathbb{N}_{0}$ which assigns to each subset $S$ of items, called a bundle, a numerical value expressing how satisfied $i$ is when $S$ is allocated to them. Additionally, there is a social impact function ${s}_{i}\colon 2^{M}\to\mathbb{N}_{0}$} for every agent $i\in{{N}}$ expressing what social impact agent $i$ generates for the society with each bundle $S\subseteq{M}$. Throughout the paper, we assume that both the valuation functions and the social impact functions are additive, meaning that for every agent $i\in{{N}}$ and every $S\subseteq{M}$ we have ${v}_{i}(S)=\sum_{g\in S}{v}_{i}(\{g\})$ and ${s}_{i}(S)=\sum_{g\in S}{s}_{i}(\{g\})$, respectively, and normalized, i.e., ${v}_{i}(\emptyset)={s}_{i}(\emptyset)=0$. An instance of fair division with social impact is then a quadruple $\mathcal{I}=({{N}},{M},({v}_{i})_{i\in{{N}}},({s}_{i})_{i\in{{N}}})$.
@@
A partial allocation is a tuple ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$, where ${A}_{i}\subseteq M$ is a (potentially empty) bundle received by agent $i\in{{N}}$, for which it holds that ${A}_{i}\cap{A}_{j}=\emptyset$ for each pair of distinct $i,j\in{{N}}$. If additionally $\bigcup_{i=1}^{{n}}{A}_{i}={M}$, then we say $\mathcal{A}$} is a complete allocation. In the remainder of the paper, we are interested in complete allocations. Let $\mathcal{A}$} be a (partial) allocation. We naturally extend the valuations and social impact functions from bundles to allocations by setting ${v}_{i}({\mathcal{A}})={v}_{i}({A}_{i})$ and ${s}_{i}({\mathcal{A}})={s}_{i}({A}_{i})$. Moreover, we denote the overall social impact generated by $\mathcal{A}$} as $\operatorname{SI}({\mathcal{A}})=\sum_{i\in{{N}}}s_{i}({A}_{i}).$
Solution Concepts.
@@
To overcome this issue, we combine SIM requirement with various notions of fairness. We focus on axioms based on pairwise comparison of agents’ bundles. The prototypical concept in this line of research is envy-freeness (Foley 1967), which requires that each agent prefers their own bundle before the bundles of other agents and is defined as follows.
@@
An allocation $\mathcal{A}$} is called envy-free (EF) if for any pair of agents $i,j\in{{N}}$ we have ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j})$.
@@
We start with a relaxation called envy-freeness up to one item (Lipton et al. 2004; Budish 2011) which requires that if there is envy from agent $i$ to agent $j$, then this envy can be eliminated by the removal of one item from $j$’s bundle.
@@
An allocation $\mathcal{A}$} is called envy-free up to one item (EF1) if for every pair of agents $i,j\in{{N}}$ there exists an item $g\in{M}$ such that ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j}\setminus\{g\})$.
Note that in the definition of EF1, the item removed from $j$’s bundle can be specific for each agent $i\in{{N}}$. A stronger variant of this notion requires that the removal of one universal item from ${A}_{j}$ prevents potential envy from all other agents. This notion is due to Conitzer et al. (2019).
@@
An allocation $\mathcal{A}$} is called strongly envy-free up to one item (sEF1) if for every $j\in{{N}}$ there exists an item $g\in{M}$ such that ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j}\setminus\{g\})$ for every $i\in{{N}}$.
As noted by Wu, Zhang, and Zhou (2025), in traditional EF1 (or sEF1), it is assumed that all agents have equal obligations. However, this assumption can be very unrealistic in many scenarios. E.g., recall our scientific laboratory example from the beginning of the paper. It is reasonable to expect that the division of the equipment should also consider factors such as seniority or merit. Therefore, we also study weighted envy-freeness (Chakraborty et al. 2021; Aziz et al. 2024; Suksompong 2025). In this model, each agent $i\in{{N}}$ is additionally associated with its weight $w_{i}$.
@@
An allocation $\mathcal{A}$} is called weighted envy-free up to one item (wEF1) if for every pair of agents $i,j\in{{N}}$ there exists an item $g\in{M}$ such that
@@
A different point of criticism towards EF1 is that it allows the envy between $i$ and $j$ to vanish simply by the removal of the best good from the $j$’s bundle, regardless of how valued this item is for the agent $i$. For example, assume that $i$ and $j$ are allocated one item each, an $i$ values ${A}_{j}$ as $1000$ and ${A}_{i}$ as $1$. Although formally there is no EF1-envy from $i$ towards $j$, agent $i$ may still consider the allocation very unfair, since there is a huge difference in the valuations of both bundles. Motivated by this, Barman et al. (2018) introduced envy-freeness up to one less preferred good.
@@
An allocation $\mathcal{A}$} is called envy-free up to one less preferred item (EFL) if for every pair of agents $i,j\in{{N}}$ at least one of the following two conditions hold:
  1. 1.
@@
An allocation $\mathcal{A}$} is called envy-free up to one transfer (tEF1) if for every pair of agents $i,j\in{{N}}$ such that ${v}_{i}({A}_{i})<{v}_{i}({A}_{j})$ there exists an item $g\in{A}_{j}$ so that ${v}_{i}({A}_{i}\cup\{g\})\geq{v}_{j}({A}_{j}\setminus\{g\})$.
@@
It is not hard to see that, given an allocation $\mathcal{A}$}, one can check whether this allocation is SIM and $\mathcal{F}$-fair, where $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$, in polynomial time. Hence, deciding whether a SIM and $\mathcal{F}$-fair allocation exists is trivially in NP and we will not stress this explicitly in our hardness proofs.
3 Socially Unaware Agents
In this section, we explore the computational complexity of deciding the existence of SIM and fair allocations. In general, hardness for EF1 follows from the result of Aziz et al. (2023a), who showed that deciding whether an EF1 allocation maximizing social-welfare [^1] Social welfare of an allocation is simply the sum of utilities of all agents in this allocation. exists is NP-complete. This corresponds to the case with ${s}_{i}={v}_{i}$ for every $i\in{{N}}$. We extend their result by studying also additional notions of fairness and giving a much more detailed picture of its complexity, even in as restricted domains as binary valuations, or with a constant number of agents.

### v2

@@ Abstract.
copyright: cccopyright: nonedoi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 6publicationmonth: 8journalyear: 2025
1. Introduction
@@
The scenarios above might be seen as fair division with externalities (velez2016fairness, 48, 12, 27), where agents care not only for the items assigned to them, but also how exactly the remaining items are allocated. Nevertheless, this model is too general, so (FlamminiGV2025, 29) have recently introduced a model that captures exactly the setting above: there is a set of goods and every agent is associated with two additive functions that define their value and social impact for each good. The goal is to allocate the goods such that the social impact is maximized while maintaining some fairness criterion. They have studied the “price of fairness”, which essentially quantifies how much worse in terms of social welfare a fair outcome is. In addition, they have formally introduced the notion of socially aware agents and which outcomes they perceive as fair. They proved that with socially aware agents, EF1 and social impact maximizing allocations always exist and can be computed in polynomial time. However, the complexity of the underlying computational problems was not examined in detail. Our goal is to extend the fairness solution concepts for this model and settle their complexity.
1.1. Our Contribution
@@
Socially-Aware Agents All Agents Socially-Aware yes NP-hard [Thm. 2] no $\alpha$-Approx. Socially-Aware yes NP-hard [Thm. 7] $\alpha<1$ Weakly Socially-Aware $\alpha=1$ NP-hard [Thm. 11] yes poly-time [Thms. 3, 5] no Binary Valuations no Valuation = Social-Impact yes poly-time [Thm. 3] yes NP-hard [Thm. 2] no Constant # of Agents no NP-hard [Thm. 2] no Bounded Valuations yes NP-hard [Thm. 4] no poly-time [Thm. 5] yes
@@
Figure 1. A basic overview of the complexity landscape of deciding whether a SIM and $\mathcal{F}$-fair allocation exist.:
@@
We provide a comprehensive study of the complexity of computing fair allocations that maximize social impact, which we term SIM allocations. We investigate seven different fairness criteria, all of which are strengthenings or weakenings of EF1. For each criterion, we examine the constraints on the valuation and social impact functions that allow for tractability (see Figure 1 for an overview of our results).
@@
Our initial set of technical results considers socially unaware agents, where we study two different dimensions of the problem and for each of them we provide complementary results between tractability and NP-hardness. Interestingly, we show that the problem behaves in the same way for every fairness criterion, albeit we need slight modifications in our proofs to formally establish this. First, we restrict the domain of the valuation and social impact functions. There, we show that the problem is NP-hard even when both functions are binary (Theorem 2), but it becomes tractable if for every agent both functions are binary and equal to each other (Theorem 3).
Then, we restrict the number of agents. Unfortunately, the problem is NP-hard even for two agents with binary social impact functions (Theorem 4). On the other hand, we derive a pseudo-polynomial time algorithm that solves the problem for any constant number of agents (Theorem 5). Hence, the problem is tractable when we have a constant number of agents, polynomially bounded valuation functions, and unrestricted social impact functions.
@@
Then we consider the case where the agents are socially aware. Our first set of results for this setting adopts the definition of social awareness from (FlamminiGV2025, 29). Intuitively, we accept allocations that do not satisfy the “standard” fairness constraint, however, any bundle an agent envies (according to the chosen fairness criterion) cannot produce strictly more social impact if this agent gets it. We prove that social awareness allows for SIM and fair allocations that can be computed in polynomial time, for all fairness criteria we study (Theorems 3 and 5). This leads us to the question of whether the definition of social awareness allows for an arbitrary fairness criterion. We prove that this is not the case by showing NP-hardness if we have “fully envious” agents (Theorem 8).
@@
Finally, we ask whether we can relax social awareness and still get a fair and SIM allocation in polynomial time. Our first relaxation considers the case where not all agents are socially aware. Unfortunately, even if only one agent is socially unaware existence of SIM and fair allocations is not guaranteed and, in fact, the problem becomes NP-hard even when there are two agents (Theorem 2). On the positive side, we complement this negative result with a pseudo-polynomial time algorithm that works for any constant number of agents where some of them are socially unaware (Theorem 4). Then, we limit social awareness for each agent in two different ways. In the first case, we focus on  $\alpha$-approximately socially aware agents, where $\alpha\in[0,1]$, and an agent accepts an unfair outcome only if every envied bundle produces at least $\alpha$ times more social impact under the current allocation. Strikingly, we get the same behavior as before: there is not always a solution and the problem is NP-hard for every possible $\alpha$ (Theorem 7). In the second case, we define weakly socially aware agents. Intuitively, an agent accepts an unfair allocation if the proportional gain of their valuation from an envied bundle is less than the proportional gain in social impact (Definition 9). Once again, there is no guaranteed solution and the problem is NP-hard even for two agents (Theorem 11).
1.2. Related Work
@@
Our paper lies in the area of fair division with indivisible items, where there is a flourish of results in the last 15 years (LiptonMMS2004, 41, 15, 18, 20); for excellent recent surveys, see (AmanatidisABFLMVW2023, 3) and (NguyenR2023, 42).
@@
As we have mentioned before, (FlamminiGV2025, 29) is the most relevant paper to ours, since we extend their model and their computational results. Having said this, both papers fall into the larger, more general model of fair division with externalities (velez2016fairness, 48, 12, 27). In this model, each agent gets value from the whole allocation, i.e., they value each bundle depending on the agent who got it. Hence, we can see any instance in our setting as an instance of agents with externalities. For socially unaware agents, we create an instance with externalities with one extra agent that has zero value for every item they get, while for socially aware agents, each agent has externalities as well that match the social impact. On a similar note, (BuLLST2023, 17) studied the model in which there is an allocator with its own view of how the items should be allocated between agents, and the goal is to be fair to both the agents and the allocator. It should be highlighted that users deriving their utility not only from their own benefit but also from the benefit of (or to) others is a well-described phenomenon in social sciences and behavioral economics; see, e.g., (Overvold1980, 43, 4, 38, 47).
@@
A different point of view is to consider our problem as constrained fair division (suksompong2021constraints, 45). We identify two settings that are closest to ours. The first regards welfare-maximizing fair allocations. Observe that when social impact equals the valuation of each agent, fair SIM allocations coincide with welfare-maximizing allocations that are fair, which was shown to be NP-hard (AzizHMS2023, 9) in general. The second is related to orientations (ChristodoulouFKS_orientations, 22, 52, 26, 1, 14). Here, every agent is associated with a subset of items that they are interested in. The goal is to find an allocation where each agent gets items only from their subset. In our case, agents are only allowed to get items for which they are social impact maximizers. The crucial difference though is that in our case, agents value the remaining items as well.
2. Preliminaries
@@
We assume a set M{M} of m{m} items which need to be partitioned between a set N{{N}} of n{{n}} agents. Each agent $i\in{{N}}$ is associated with a valuation function ${v}_{i}\colon 2^{M}\to\mathbb{N}_{0}$ which assigns to each subset $S$ of items, called a bundle, a numerical value expressing how satisfied $i$ is when $S$ is allocated to them. Additionally, there is a social impact function ${s}_{i}\colon 2^{M}\to\mathbb{N}_{0}$} for every agent $i\in{{N}}$ expressing what social impact agent $i$ generates for the society with each bundle $S\subseteq{M}$. Throughout the paper, we assume that both the valuation functions and the social impact functions are additive, meaning that for every agent $i\in{{N}}$ and every $S\subseteq{M}$ we have ${v}_{i}(S)=\sum_{g\in S}{v}_{i}(\{g\})$ and ${s}_{i}(S)=\sum_{g\in S}{s}_{i}(\{g\})$, respectively, and normalized, i.e., ${v}_{i}(\emptyset)={s}_{i}(\emptyset)=0$. If an item $g\in{M}$ is valued (in terms of $v_{i}$’s) non-negatively by every agent, we say that $g$ is good. An item is called a chore if it is valued non-positively by every agent. Unless stated otherwise, we assume that all items are goods. An instance of fair division with social impact is then a quadruple $\mathcal{I}=({{N}},{M},({v}_{i})_{i\in{{N}}},({s}_{i})_{i\in{{N}}})$.
@@
A partial allocation is a tuple ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$, where ${A}_{i}\subseteq M$ is a (potentially empty) bundle received by agent $i\in{{N}}$, for which it holds that ${A}_{i}\cap{A}_{j}=\emptyset$ for each pair of distinct $i,j\in{{N}}$. If additionally $\bigcup_{i=1}^{{n}}{A}_{i}={M}$, then we say $\mathcal{A}$} is a complete allocation. In the remainder of the paper, we are interested in complete allocations. Let $\mathcal{A}$} be a (partial) allocation. We naturally extend the valuations and social impact functions from bundles to allocations by setting ${v}_{i}({\mathcal{A}})={v}_{i}({A}_{i})$ and ${s}_{i}({\mathcal{A}})={s}_{i}({A}_{i})$. Moreover, we denote the overall social impact generated by $\mathcal{A}$} as
$\operatorname{SI}({\mathcal{A}})=\sum_{i\in{{N}}}s_{i}({A}_{i}).$
2.1. Solution Concepts
@@
To overcome this issue, we combine SIM requirement with various notions of fairness. We focus on axioms based on pairwise comparison of agents’ bundles. The prototypical concept in this line of research is envy-freeness (Foley1967, 30), which requires that each agent prefers their own bundle before the bundles of other agents and is defined as follows.
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called envy-free (EF) if for any pair of agents $i,j\in{{N}}$ we have ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j})$.
@@
We start with a relaxation called envy-freeness up to one item (LiptonMMS2004, 41, 18) which requires that if there is envy from agent $i$ to agent $j$, then this envy can be eliminated by the removal of one item from $j$’s bundle.
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called envy-free up to one item (EF1) if for every pair of agents $i,j\in{{N}}$ there exists an item $g\in{M}$ such that ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j}\setminus\{g\})$.
Note that in the definition of EF1, the item removed from $j$’s bundle can be specific for each agent $i\in{{N}}$. A stronger variant of this notion requires that the removal of one universal item from ${A}_{j}$ prevents potential envy from all other agents. This notion is due to (ConitzerFSW2019, 23).
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called strongly envy-free up to one item (sEF1) if for every $j\in{{N}}$ there exists an item $g\in{M}$ such that ${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j}\setminus\{g\})$ for every $i\in{{N}}$.
As noted by (WuZZ2025, 50), in traditional EF1 (or sEF1), it is assumed that all agents have equal obligations. However, this assumption can be very unrealistic in many scenarios. E.g., recall our scientific laboratory example from the beginning of the paper. It is reasonable to expect that the division of the equipment should also consider factors such as seniority or merit. Therefore, we also study weighted envy-freeness (ChakrabortyISZ2021, 21, 10, 46). In this model, each agent $i\in{{N}}$ is additionally associated with its weight $w_{i}$. The concept is then defined as follows.
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called weighted envy-free up to one item (wEF1) if for every pair of agents $i,j\in{{N}}$ there exists an item $g\in{M}$ such that
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called strongly weighted envy-free up to one item (swEF1) if for every $j\in{{N}}$ there exists an item $g\in{M}$ such that for every $i\in{{N}}$ we have
$\frac{{v}_{i}({A}_{i})}{w_{i}}\geq\frac{{v}({A}_{j}\setminus\{g\})}{w_{j}}\,.$
A different point of criticism towards EF1 is that it allows the envy between $i$ and $j$ to vanish simply by the removal of the best good from the $j$’s bundle, regardless of how valued this item is for the agent $i$. For example, assume that $i$ and $j$ are allocated one item each, an $i$ values ${A}_{j}$ as $1000$ and ${A}_{i}$ as $1$. Although formally there is no EF1-envy from $i$ towards $j$, agent $i$ may still consider the allocation very unfair, since there is a huge difference in the valuations of both bundles. Motivated by this, (BarmanBMN2018, 13) introduced envy-freeness up to one less preferred good.
Definition 0.
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called envy-free up to one less preferred item (EFL) if for every pair of agents $i,j\in{{N}}$ at least one of the following two conditions hold:
  1. (1)
@@
It follows directly from the definition that every EFL allocation is simultaneously EF1.
@@
An allocation ${\mathcal{A}}=({A}_{1},\ldots,{A}_{{n}})$ is called envy-free up to one transfer (tEF1) if for every pair of agents $i,j\in{{N}}$ such that ${v}_{i}({A}_{i})<{v}_{i}({A}_{j})$ there exists an item $g\in{A}_{j}$ so that ${v}_{i}({A}_{i}\cup\{g\})\geq{v}_{j}({A}_{j}\setminus\{g\})$.
@@
It is not hard to see that given an allocation $\mathcal{A}$}, one can check whether this allocation is SIM and $\mathcal{F}$-fair, where $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$, in polynomial time. Hence, deciding whether a SIM and $\mathcal{F}$-fair allocation exist is trivially in NP and we will not stress this explicitly in our hardness proofs.
3. Socially Unaware Agents
In this section, we explore the computational complexity of deciding the existence of SIM and fair allocations. In general, hardness for EF1 follows from the result of (AzizHMS2023, 9), who showed that deciding whether an EF1 allocation maximizing social-welfare [^1] Social welfare of an allocation is simply the sum of utilities of all agents in this allocation. exists is NP-complete. This corresponds to the case with ${s}_{i}={v}_{i}$ for every $i\in{{N}}$. We extend their result by studying also additional notions of fairness and giving a much more detailed picture of its complexity, even in as restricted domains as binary valuations, or with a constant number of agents.